I'm trying to find n in the following:
1000 * sin(n/2) * 1.1 = 285.5
where / is divided by.
Thank you.

Answer by Alan3354(69443)

I'm trying to find n in the following:
1000 * sin(n/2) * 1.1 = 285.5
-----------------
sin(n/2) = 285.5/(1.1*1000)